Hello. I ran across a problem that left me baffled. It is not a hard one but I just did not retain it whe nI learned algebra if I learned it at all. Here goes:
The constraint 5x+2y<=0 passes through (20,50), (10,0), (20,0), (0,-5). Why and how?
Also, the constraint 2x-3y>=0 passes through (10,50), (30, 20), (0, 30) (-10, 0)
Why and how?

Answer by stanbon(75887)

The constraint 5x+2y<=0 passes through (20,50), (10,0), (20,0), (0,-5). Why and how?
Substitute x and y in the inequality with the numbers in each numbered pair
to see if you get a true statement.
Example: using (20,50) you get
5*20+2*50 <= 0
100 + 100 <= 0
200 <= 0
That is false.
---------------
It is true only for (0,-5)
------------------------
Also, the constraint 2x-3y>=0 passes through (10,50), (30, 20), (0, 30) (-10, 0)
Why and how?
==============================
Cheers,
Stan H.